Tax Savings and Costing (The Case of Transfer Pricing)

Please prepare a report answering the listed questions. You may Excel to create spreadsheets and copy the answers to this document.

Hansen, Kotter, and Zales is a law firm that contains one service department (Research & Document) and two production departments (Litigation and Consulting). The firm employs a job-order costing system to accumulate costs chargeable to each client. The firm uses actual costing to assign overhead. General overhead costs can be allocated based on either direct attorney hours or the number of employees, depending on managers’ choice. At the end of the year, the records revealed the actual general overhead costs are $720,000. At the end of the year, the records revealed the following costs and operating data for all cases handled during the year:

Research & Document

Litigation

Consulting

# of Employees

10

8

6

Direct Attorney Hours(# of hrs)

3,000

8,000

5,000

Direct Attorney Costs ($)

$150,000

$400,000

$250,000

Direct Material Costs ($)

$16,000

$15,500

$13,500

*** 50% of Research & Department's service is provided to litigation department and the other 50% to consulting department.

Part I Cost Allocation

A. Compute the overhead allocation rates for general overhead based on different cost drivers. What are the overhead costs assigned to each department, using different cost drivers?

B. Compare the total costs of each production department after departmental cost allocation, using attorney hours and the number of employees as cost driver, respectively. Does the choice of cost driver affect the total costs of each production department?
Answer #1
OVERHEAD COST ALLOCATION BASED ON DIRECT ATTORNEY HOUR:
Research & Development Litigation Consulting Total
A Direct Attorney Hours 3000 8000 5000 16000
B=A/16000 Percentage of total 18.75% 50.00% 31.25% 100.00%
C=B*720000 Allocated Overhead Cost $135,000 $360,000 $225,000 $720,000
OVERHEAD COST ALLOCATION BASED ON NUMBER OF EMPLOYEES:
Research & Development Litigation Consulting Total
A Number of Employees 10 8 6 24
B=A/24 Percentage of total 41.67% 33.33% 25.00% 100.00%
C=B*720000 Allocated Overhead Cost $300,000 $240,000 $180,000 $720,000
B OVERHEAD COST OF PRODUCTION DEPARTMENTS
BASED ON ATTORNEY HOURS
Litigation Consulting Total
Allocated Costs $360,000 $225,000 $585,000
Research & Development Costs apprtioned $67,500 $67,500 $135,000 (50%*135000)
Total Overhead Costs $427,500 $292,500 $720,000
OVERHEAD COST OF PRODUCTION DEPARTMENTS
BASED ON NUMBER OF EMPLOYEES
Litigation Consulting Total
Allocated Costs $240,000 $180,000 $420,000
Research & Development Costs apprtioned $150,000 $150,000 $300,000 (50%*300000)
Total Overhead Costs $390,000 $330,000 $720,000
YES, Choice of cost driver affects cost of each production department